# EARN Fairness: Explaining, Asking, Reviewing, and Negotiating Artificial Intelligence Fairness Metrics Among Stakeholders

### v2

@@ 6. Discussion
In this section, we will first discuss the limitations of our work and future research. Then, we will discuss implications for the design of fairness tools and processes by considering how UIs could be extended to express personal preferences, the complexity of accommodating diverse personal fairness preferences, how to support negotiation in groups, and how to alleviate problems in reaching consensus. We reflect on how these tools and processes would need to scale up to take into account larger and distributed stakeholder groups.
@@ 6.2. Implications for Human-Centered Fairness
We argue that tackling AI fairness transcends the bounds of solely technical, AI expert-driven approaches, evolving into a broader, human-centered concern. Our work demonstrates the feasibility of this approach. We will discuss how our work presenting the EARN Fairness framework and the FEE interactive systems has implications for design to support stakeholders’ elicitation of personal preferences and negotiation of consensus fairness metrics.
@@ 6.2.1. Explaining Fairness Metrics
Explaining fairness metrics through FEE also had the effect of engaging the participants in our study to think about fairness, how fairness is defined, and what fairness means to them. For example, participants excluded metrics that defined fairness within specific groups, such as Equal Opportunity, which only ensured fairness within the subgroup that had been labeled as ”Good Credit”. Furthermore, we observed that when participants made modifications to AI predictions or ground-truth labels to explore real-time updates in fairness results, this made them ”reflect or re-think how the fairness results changed” (P14), enhancing their understanding of the metrics. Therefore, we believe that soliciting stakeholders’ metric preferences based on exposing the underlying logic offers a robust approach.
@@ 6.2.2. Complexity of Diverse Fairness Preferences
This leads to the question of what to do with different personal preferences once we have elicited them, in order to accommodate differences. Given that there is no one-size-fits-all metric and selecting only one metric is insufficient (Mehrabi et al., 2021), one possible approach for AI practitioners and researchers, as revealed through our findings, is preference integration. When teams negotiated their preferred fairness metric, we observed a common pattern: all teams tended to merge diverse preferences to form a more comprehensive and inclusive definition of AI fairness to resolve differences, despite using different negotiation strategies. That is, in the credit rating scenario, all teams agreed on a dual focus on the absence of prejudice or favoritism toward both individuals and groups. While individual and group fairness might not inherently conflict as they reflect different aspects of consistent moral and political concerns (Binns, 2020), conflicts may arise in practice when optimizing specific metrics (Mehrabi et al., 2021). We will next discuss how negotiations around conflicting personal preferences could be supported.
@@ 6.2.5. Scaling Up The Process
To support the negotiation among distributed groups of people, developing an online system for a larger group is essential. This could be achieved by integrating synchronous and asynchronous online collaboration tools. For example, other tools (Zhang et al., 2023; Kawakami et al., 2024) allow personal preference data to be imported into online whiteboards to display the results to a group of stakeholders and attach reasoning or justifications, thus providing a foundation for deliberation. However, these tools will not be adequate for larger groups as negotiation support may quickly become unwieldy with too many participants. How to address this challenge is an open question but could be solved with other democratic processes already in place, for example, by choosing stakeholder representatives.
@@ 7. Conclusions
  • Lay stakeholders differ from popular metrics in industry and academia, and there is no overall preference for one metric. There is a need for consensus to be negotiated.
@@
  • We observed some problems in reaching consensus during the negotiations, mainly centering on ways to resolve conflicts. For consensus, groups typically chose to compromise by combining multiple metrics to address differences in personal preferences.
We discussed the implications for future tools and processes that support reaching consensus on choosing fairness metrics. Our framework and interactive system present a feasible step for engaging stakeholders in the expression of personal preferences and collective selection of metrics, fostering more equitable and inclusive AI fairness.
@@ A.1. Fairness Metrics and Explanations
Group Fairness1 Demographic Parity (Kusner et al., 2017), (Dwork et al., 2012) Both protected and unprotected group members have an equal likelihood of being classified in the positive category. Difference = $P(\hat{Y}=\text{1}\mid\text{G = 0})-P(\hat{Y}=\text{1}\mid\text{G = 1})$
@@
Individual Fairness Counterfactual Fairness (Kusner et al., 2017) An individual receives the same decision in both the actual world and a counterfactual world where the individual belonged to a different demographic group. $CFR=\frac{1}{N}\sum_{(X,A_{\text{true}})\in\text{dataset}}1\{p(X,A_{\text{true}})=p(X,A_{\text{counterfactual}})\}$ 2
@@ A.3.2. Participants’ Reasons for Selection and Exclusion
Consistency Reasons for Selection Comparison with Similar Individuals Expressing a desire to compare their results with those of similar individuals to ensure fairness. If people similar to me predictably have good credit, I should also receive a favorable prediction. It’s fair. It allows you to compare yourself with people of the same level or qualifications. (P14) 6 7
Inclusivity for Individuals Expressing a preference for its potential to prevent discrimination and ensure no one is unfairly excluded. ’Cause it prevents discrimination like, no one’s getting left out sort of thing [among similar individuals]. (P13) 2 2
Decision Transparency Expressing a desire to compare their situation with others as a means to understand their own results. I can compare to others to understand why I cannot achieve better results. (P12) 1 1
Reasons for Exclusion Similarity Is Quantified Expressing users’ concern about the inherent subjectivity in defining and measuring similarity, and the difficulty to objectively quantify. There is no way or hard to find a standard way to analyze and compare similarity objectively. (P16) 5 8
Counterfactual Fairness Reasons for Selection Irrelevance Of Protected Features Expressing a belief that factors traditionally associated with bias, such as age, gender, or foreign worker status, i.e., protected features, should not influence AI’s prediction results. Because it disregards their age, their gender and their foreign worker status. […] Nothing should affect it at all. (P17) 5 6
Individual Characteristic-based Consideration Expressing a preference for a detailed, targeted consideration of how characteristic (i.e., protected features) would affect an individual, as opposed to broader, more generalized assessments. The factors that are that mostly contribute to a bias […] you’re looking at it on an individual basis. […] You take their individual characteristics into account. (P10) 4 7
Objective Qualification Expressing a preference for counterfactual fairness because it offers a more objective and convincing way to quantify individual fairness. It is like I can quantity the similarity because the overall information for this person is still not changed, we just change the personal characteristics, [… which] is more convincing. (P5) 3 5
Reasons for Exclusion Practical Ineffectiveness Expressing users’ doubts about the effectiveness of fairness evaluation, driven by the belief that altering a single feature value inherently wouldn’t impact AI predictions. It’s not that useless. But It’s not that helpful. Because the digital I think they are mostly the same, and however, I don’t think there are so many big, big differences. (P2) 1 2
Equalized Odds Reasons for Selection Disparity Control for all Parties Expressing the users’ preference for a approach to control disparities between different groups. we need to protect those vulnerable people, […] [rated] bad [credit] people. (P14) 4 5
Integrated Fairness metric Expressing users’ preference for Equalized Odds as it integrates different fairness metrics to help detect various bias types. Like combine them [Equal Opportunity and Predictive equality] together, you’re taking into account both biases that could take place here. This makes more sense. (P10) 3 4
Comprehensive Complaint Handling Expressing the users’ preference for Equalized Odds due to its wide-ranging and thorough approach to managing and resolving complaints from all customer groups. Although my targets are not bad people, I will choose Equalized Odds, because, we have coping strategies for each kind of customer.(P16) 2 2
Ground-Truth Label Importance Expressing user’s preference for metrics that rely on rated credits and belief that such alignment ensures a more trustworthy assessment of fairness. I think the rated credits will be more believable part to evaluate this because. I think it’s mostly come from the rated one.[…] I really think the rated credit is more important than the condition features. (P2) 1 2
Reasons for Exclusion Redundancy Expressing that using Equalized Odds introduces a redundancy of fairness considerations, when users think pursuing fairness should be sufficient for qualified people. I think it’s good just to focus on the good part. (P4) 1 1
Equal Opportunity Reasons for Selection Qualified People Focus Only Expressing a preference for only focusing on individuals with good qualifications when making decisions I think in the loan applications, we have to care about if the customer has good qualifications. If the customer has already been judged as bad credit by humans, we do not need to care about these sections at all. (P4) 5 11
Reasons for Exclusion Incompleteness Expressing that Equal Opportunity only assesses fairness among qualified people, overlooking comprehensive fairness when users prefer fairness for all, regardless of qualifications. I think all the people are important. The overall, the whole. (P6) 1 1
Predictive Equality Reasons for Selection Risk Considerations on Non-Qualified People Expressing a perspective focused on risk management and ethical considerations which suggests a belief in the need for both accurate and fair decisions. Like bad people should not be provided with loans. […]. We should provide more care in making sure that bad people don’t get loads. (P15) 2 4
Reasons for Exclusion Unnecessary to consider non-qualified individuals Expressing that using Predictive Equality considering fairness among non-qualified people is unnecessary and meaningless, when users think pursuing fairness is only for qualified people. I will not consider the fairness among bad people. Meaningless. For the loan application, […] the final purpose is to provide a better service to those qualified, potential customers. (P16) 4 6
Outcome Test Reasons for Selection Direct Fairness Assessment Expressing the users’ belief that the fairness of a model can be assessed by directly comparing the model’s predictions with actual labels. With good predictions and bad predictions, we are trying to figure out which are the good which are the bad. And based on that, we are trying to […] get [fair] predictions. I feel like that’s the case. P(18) 3 4
Reasons for Exclusion Over-Reliance on AI’s Performance Expressing unease with first filtering out the AI’s ’good’ predictions and then assessing the proportion of real good ones, especially when doubting AI’s effectiveness. Because the outcome test is pretty rely on the AI’s predictions. our current model accuracy is only 76%. (P9) 2 2
@@
Demographic Parity Reasons for Selection Holistic Fairness Assessment Expressing a preference for considering a holistic assessment where the total population of each group is accounted for, rather than focusing on further refined groups based on types of ground-truth labels or prediction results. The number of people you judged or cared about is more comprehensive. If you just focused on like rated good or rated bad […] you always divided them into two. But demographic parity could give me an overall percentage in all group people. (P5) 2 2
Reasons for Exclusion Neglect Individual Qualifications Expressing users’ concerns about the neglect of people’s qualifications in the pursuit of fairness. It doesn’t actually care about real good or bad. It just gives importance to the gender equality thing. (P15) 5 7
@@ A.3.3. Participants’ Feedback on UI Design
Table 8. Codebook
@@
Suggestions for Improving UI Design Model Transparency Expressing that users engage with AI predictions, they seek to identify the specific factors that have a greater impact on the final results. But the person reviewing them might also want to look at the reason for that. Maybe you could […] give me what feature was it about the customer that caused the decision to be bad. (P13) 2 3
Practical Implications of Fairness Metrics Expressing that users’ need for both technical knowledge of metric calculations and insights into their practical applications and real-world implications, addressing difficulties in metric interpretation due to abstract naming and a lack of contextual clarity. If each metric combining with a context explanation would be better for me to select. (P16) I can understand how they are calculated but don’t know what the point is. (P3) 6 10
Simplified Data Visualization & Explanation Expressing the need to tailor data visualization formats to user expertise, starting with basic, universally accessible formats like bar and pie charts for conveying fairness concepts to lay users. Maybe also like bar charts are just easier to read for everybody. So expressed in more like pie chart formats or easier formats or like regular bar graphs, that might be easier for like fully lay people to understand if that makes sense. (P10) 3 3
Simplifying Information Quantity Expressing the participant’s viewpoint on limiting the number of data points visible to users at any given time, aiming to alleviate cognitive overload and prevent confusion that might arise from an abundance of data. At least at least for the UI initially like maybe have less data points visible […] why is there so much data? What am I supposed to get gather from that? […] let’s say it might be more difficult for them to gather anything. (P10) 3 4
User Experience Cognitively Challenging Expressing a user experience that requires a high level of mental effort and concentration. I got a little bit lost at first so until I finish the whole process. […] It makes a lot of time to understand. Get to half an hour for me. (P8) 8 12
Engaging Expressing users find this UI prototype and the AI fairness content engaging and interesting. I think it is very interesting for me. I can understand your research. (P6) 10 10

### v3

@@ 6. Discussion
In this section, we will first discuss the limitations of our work and future research. Then, we will discuss the implications for the design of fairness tools and processes by considering how UIs could be extended to express personal preferences, the complexity of accommodating diverse personal fairness preferences, how to support negotiation in groups, and how to alleviate problems in reaching consensus. We reflect on how these tools and processes would need to scale up to take into account larger and distributed stakeholder groups.
@@ 6.2. Implications for Human-Centered Fairness
We argue that tackling AI fairness transcends the bounds of solely technical, AI expert-driven approaches, evolving into a broader, human-centered concern. Our work demonstrates the feasibility of this approach. We will discuss how our work presenting the EARN Fairness framework and the FEE interactive system has implications for design to support stakeholders’ elicitation of personal preferences and negotiation toward fairness metric consensus.
@@ 6.2.1. Explaining Fairness Metrics
Explaining fairness metrics through FEE also had the effect of engaging the participants in our study to think about fairness, how fairness is defined, and what fairness means to them. For example, participants excluded metrics that define fairness within specific groups, such as Equal Opportunity, which ensures fairness only within the subgroup labeled ”Good Credit”. Furthermore, we observed that when participants made modifications to AI predictions or ground-truth labels to explore real-time updates in fairness results, it prompted them ”reflect or re-think how the fairness results changed” (P14), enhancing their understanding of the metrics. Therefore, we believe that soliciting stakeholders’ metric preferences based on exposing the underlying logic offers a robust approach.
@@ 6.2.2. Complexity of Diverse Fairness Preferences
This leads to the question of what to do with different personal preferences once we have elicited them, in order to accommodate differences. Given that there is no ”one-size-fits-all” metric and selecting only one metric is insufficient (Mehrabi et al., 2021), one possible approach for AI practitioners and researchers, as revealed through our findings, is preference integration. When teams negotiated their preferred fairness metric, we observed a common pattern: all teams tended to merge diverse preferences to form a more comprehensive and inclusive definition of AI fairness to resolve differences, despite using different negotiation strategies. That is, in the credit rating scenario, all teams agreed on a dual focus on the absence of prejudice or favoritism toward both individuals and groups. While individual and group fairness might not inherently conflict as they reflect different aspects of consistent moral and political concerns (Binns, 2020), conflicts may arise in practice when optimizing specific metrics (Mehrabi et al., 2021). We will next discuss how negotiations around conflicting personal preferences could be supported.
@@ 6.2.5. Scaling Up the Process
To support negotiation among distributed groups of people, developing an online system for a larger group is essential. This could be achieved by integrating synchronous and asynchronous online collaboration tools. For example, other tools (Zhang et al., 2023; Kawakami et al., 2024) allow personal preference data to be imported into online whiteboards to display the results to a group of stakeholders and attach reasoning or justifications, thus providing a foundation for deliberation. However, these tools may not be adequate for larger groups, as negotiation support can quickly become unwieldy with too many participants. How to address this challenge is an open question, but it could be solved with other democratic processes already in place, for example, by choosing stakeholder representatives.
@@ 7. Conclusions
  • Lay stakeholders’ metric preferences differ from popular metrics in industry and academia, and there is no overall preference for one metric. There is a need for consensus to be negotiated among stakeholders.
@@
  • We observed some problems in reaching consensus during the negotiations, mainly centering on ways to resolve conflicts. For consensus, groups typically choose to compromise by combining multiple metrics to address differences in personal preferences.
We discussed the implications for future tools and processes that support reaching consensus on choosing fairness metrics. Our framework, comprising the EARN Fairness process and the interactive system, presents a feasible step for engaging stakeholders in the expression of personal preferences and collective selection of metrics, fostering more equitable and inclusive AI fairness.
@@ A.1. Fairness Metrics and Explanations
1Group Fairness Demo-graphic Parity (Kusner et al., 2017), (Dwork et al., 2012) Both protected and unprotected group members have an equal likelihood of being classified in the positive category. Difference = $P(\hat{Y}=\text{1}\mid\text{G = 0})-P(\hat{Y}=\text{1}\mid\text{G = 1})$
@@
[…]
@@ A.3.2. Participants’ Reasons for Selection and Exclusion
Continued on next page.
Table 7 Continued from previous page Theme Subtheme Code Code Description Example Pts. Freq. Consistency Reasons for Selection Comparison with Similar Individuals Expressing a desire to compare their results with those of similar individuals to ensure fairness. If people similar to me predictably have good credit, I should also receive a favorable prediction. It’s fair. It allows you to compare yourself with people of the same level or qualifications. (P14) 6 7 Inclusivity for Individuals Expressing a preference for its potential to prevent discrimination and ensure no one is unfairly excluded. ’Cause it prevents discrimination like, no one’s getting left out sort of thing [among similar individuals]. (P13) 2 2 Decision Transparency Expressing a desire to compare their situation with others as a means to understand their own results. I can compare to others to understand why I cannot achieve better results. (P12) 1 1 Reasons for Exclusion Similarity Is Quantified Expressing users’ concern about the inherent subjectivity in defining and measuring similarity, and the difficulty to objectively quantify. There is no way or hard to find a standard way to analyze and compare similarity objectively. (P16) 5 8 Counterfactual Fairness Reasons for Selection Irrelevance Of Protected Features Expressing a belief that factors traditionally associated with bias, such as age, gender, or foreign worker status, i.e., protected features, should not influence AI’s prediction results. Because it disregards their age, their gender and their foreign worker status. […] Nothing should affect it at all. (P17) 5 6 Individual Characteristic-based Consideration Expressing a preference for a detailed, targeted consideration of how characteristic (i.e., protected features) would affect an individual, as opposed to broader, more generalized assessments. The factors that are that mostly contribute to a bias […] you’re looking at it on an individual basis. […] You take their individual characteristics into account. (P10) 4 7
Continued on next page.
Continued from previous page Theme Subtheme Code Code Description Example Pts. Freq. Objective Qualification Expressing a preference for counterfactual fairness because it offers a more objective and convincing way to quantify individual fairness. It is like I can quantity the similarity because the overall information for this person is still not changed, we just change the personal characteristics, [… which] is more convincing. (P5) 3 5 Reasons for Exclusion Practical Ineffectiveness Expressing users’ doubts about the effectiveness of fairness evaluation, driven by the belief that altering a single feature value inherently wouldn’t impact AI predictions. It’s not that useless. But It’s not that helpful. Because the digital I think they are mostly the same, and however, I don’t think there are so many big, big differences. (P2) 1 2 Equalized Odds Reasons for Selection Disparity Control for all Parties Expressing the users’ preference for a approach to control disparities between different groups. we need to protect those vulnerable people, […] [rated] bad [credit] people. (P14) 4 5 Integrated Fairness metric Expressing users’ preference for Equalized Odds as it integrates different fairness metrics to help detect various bias types. Like combine them [Equal Opportunity and Predictive equality] together, you’re taking into account both biases that could take place here. This makes more sense. (P10) 3 4 Comprehensive Complaint Handling Expressing the users’ preference for Equalized Odds due to its wide-ranging and thorough approach to managing and resolving complaints from all customer groups. Although my targets are not bad people, I will choose Equalized Odds, because, we have coping strategies for each kind of customer.(P16) 2 2 Ground-Truth Label Importance Expressing user’s preference for metrics that rely on rated credits and belief that such alignment ensures a more trustworthy assessment of fairness. I think the rated credits will be more believable part to evaluate this because. I think it’s mostly come from the rated one.[…] I really think the rated credit is more important than the condition features. (P2) 1 2
Continued on next page.
Continued from previous page Theme Subtheme Code Code Description Example Pts. Freq. Reasons for Exclusion Redundancy Expressing that using Equalized Odds introduces a redundancy of fairness considerations, when users think pursuing fairness should be sufficient for qualified people. I think it’s good just to focus on the good part. (P4) 1 1 Equal Opportunity Reasons for Selection Qualified People Focus Only Expressing a preference for only focusing on individuals with good qualifications when making decisions I think in the loan applications, we have to care about if the customer has good qualifications. If the customer has already been judged as bad credit by humans, we do not need to care about these sections at all. (P4) 5 11 Reasons for Exclusion Incompleteness Expressing that Equal Opportunity only assesses fairness among qualified people, overlooking comprehensive fairness when users prefer fairness for all, regardless of qualifications. I think all the people are important. The overall, the whole. (P6) 1 1 Predictive Equality Reasons for Selection Risk Considerations on Non-Qualified People Expressing a perspective focused on risk management and ethical considerations which suggests a belief in the need for both accurate and fair decisions. Like bad people should not be provided with loans. […]. We should provide more care in making sure that bad people don’t get loads. (P15) 2 4 Reasons for Exclusion Unnecessary to consider non-qualified individuals Expressing that using Predictive Equality considering fairness among non-qualified people is unnecessary and meaningless, when users think pursuing fairness is only for qualified people. I will not consider the fairness among bad people. Meaningless. For the loan application, […] the final purpose is to provide a better service to those qualified, potential customers. (P16) 4 6
Continued on next page.
@@
Continued from previous page Theme Subtheme Code Code Description Example Pts. Freq. Outcome Test Reasons for Selection Direct Fairness Assessment Expressing the users’ belief that the fairness of a model can be assessed by directly comparing the model’s predictions with actual labels. With good predictions and bad predictions, we are trying to figure out which are the good which are the bad. And based on that, we are trying to […] get [fair] predictions. I feel like that’s the case. P(18) 3 4 Reasons for Exclusion Over-Reliance on AI’s Performance Expressing unease with first filtering out the AI’s ’good’ predictions and then assessing the proportion of real good ones, especially when doubting AI’s effectiveness. Because the outcome test is pretty rely on the AI’s predictions. our current model accuracy is only 76%. (P9) 2 2 Demographic Parity Reasons for Selection Holistic Fairness Assessment Expressing a preference for considering a holistic assessment where the total population of each group is accounted for, rather than focusing on further refined groups based on types of ground-truth labels or prediction results. The number of people you judged or cared about is more comprehensive. If you just focused on like rated good or rated bad […] you always divided them into two. But demographic parity could give me an overall percentage in all group people. (P5) 2 2 Reasons for Exclusion Neglect Individual Qualifications Expressing users’ concerns about the neglect of people’s qualifications in the pursuit of fairness. It doesn’t actually care about real good or bad. It just gives importance to the gender equality thing. (P15) 5 7
@@ A.3.3. Participants’ Feedback on UI Design
Below is the codebook we generated while analyzing participants feedback.
Table 8. Codebook-Participants’ Feedback on UI Design
@@
Continued on next page.
Table 8 Continued from previous page Subtheme Code Code Description Example Pts. Freq. Suggestions for Improving UI Design Model Transparency Expressing that users engage with AI predictions, they seek to identify the specific factors that have a greater impact on the final results. But the person reviewing them might also want to look at the reason for that. Maybe you could […] give me what feature was it about the customer that caused the decision to be bad. (P13) 2 3 Practical Implications of Fairness Metrics Expressing that users’ need for both technical knowledge of metric calculations and insights into their practical applications and real-world implications, addressing difficulties in metric interpretation due to abstract naming and a lack of contextual clarity. If each metric combining with a context explanation would be better for me to select. (P16) I can understand how they are calculated but don’t know what the point is. (P3) 6 10 Simplified Data Visualization & Explanation Expressing the need to tailor data visualization formats to user expertise, starting with basic, universally accessible formats like bar and pie charts for conveying fairness concepts to lay users. Maybe also like bar charts are just easier to read for everybody. So expressed in more like pie chart formats or easier formats or like regular bar graphs, that might be easier for like fully lay people to understand if that makes sense. (P10) 3 3 Simplifying Information Quantity Expressing the participant’s viewpoint on limiting the number of data points visible to users at any given time, aiming to alleviate cognitive overload and prevent confusion that might arise from an abundance of data. At least at least for the UI initially like maybe have less data points visible […] why is there so much data? What am I supposed to get gather from that? […] let’s say it might be more difficult for them to gather anything. (P10) 3 4 User Experience Cognitively Challenging Expressing a user experience that requires a high level of mental effort and concentration. I got a little bit lost at first so until I finish the whole process. […] It makes a lot of time to understand. Get to half an hour for me. (P8) 8 12 Engaging Expressing users find this UI prototype and the AI fairness content engaging and interesting. I think it is very interesting for me. I can understand your research. (P6) 10 10